OpenStudy (anonymous):

Help writing the equation y - 4 = -2(x - 5) in slope-intercept form? How would I convert it? Medal and fan!!

OpenStudy (aum):

Slope-Intercept form is:\[ y = mx + b \]y - 4 = -2(x-5) distribute the right hand side: y - 4 = -2x + 10 add 4: y = -2x + 10 + 4 y = -2x + 14 This is the slope-intercept form.

OpenStudy (anonymous):

Awesome, thanks!

OpenStudy (aum):

You are welcome.